OpenStudy (anonymous):

Jim Gym, a high school basketball coach, wants to analyze the height of the boys at his school. He knows their height is normally distributed so he can use the standard normal distribution. He measures the height of 100 randomly selected boys. Next, he calculates the mean and standard deviation of their heights. The mean is 66 inches and the standard deviation is 5 inches. Jim uses the normal distribution table to calculate the number of boys in each segment of the distribution.

OpenStudy (amistre64):

Jim seems busy .....

OpenStudy (anonymous):

yes very!

OpenStudy (amistre64):

there seems to be anough information given to answer some question ... but there isnt any question posted that relates to the information

OpenStudy (anonymous):

Standard Deviation Percentage from table Number of boys out of 100 -2 to -1 (a0 to 61 inches) 14 -1 to 0 (61 to a2 inches) 34.1% 0 to +1 (66 to 71 inches) +1 to +2 (71 to a6 inches) 13.6 %

OpenStudy (anonymous):

the a0, a2, and a6 are blank spaces for filling in

OpenStudy (amistre64):

and of course they would correspond to the z values related to -2 sds, 0sds, and 2sds

OpenStudy (anonymous):

yes do you know the chart because I have it

OpenStudy (amistre64):

since the distribution is symetric about the mean, we can fill in some of this stuff -2 to -1 (a0 to 61 inches) 13.6% 14 -1 to 0 (61 to a2 inches) 34.1% n 0 to +1 (66 to 71 inches) 34.1% n +1 to +2 (71 to a6 inches) 13.6 % 14

OpenStudy (anonymous):

so how do I find the others?

OpenStudy (amistre64):

out of well, since 34.1 percent of 100 is 34.1 boys, lets say thats 35

OpenStudy (amistre64):

or should we round down to 34 since they rounded up the other 2?

OpenStudy (amistre64):

i say round up since we have the 3sds to cover any slack with

OpenStudy (anonymous):

that seems too low.

OpenStudy (anonymous):

yes round up

OpenStudy (amistre64):

-2 to -1 (a0 to 61 inches) 13.6% 14 -1 to 0 (61 to a2 inches) 34.1% 35 0 to +1 (66 to 71 inches) 34.1% 35 +1 to +2 (71 to a6 inches) 13.6 % 14

OpenStudy (amistre64):

since the sd is 5 ... that means there is a range of 5 between each .... thing they give you

OpenStudy (amistre64):

a0 + 5 = 61 61+5 = a2 66 + 5 = 71 71+5 = a6

OpenStudy (anonymous):

that makes perfect sense!!

OpenStudy (amistre64):

i know right!! :) you should be able to wrap this up now

OpenStudy (anonymous):

of course thank you very much!! i hope I do good;)